# Convergence properties of $T^{\prime}$-Expansion Scheme: Hadron Resonance Gas and Cluster Expansion Model

### v1

@@ I INTRODUCTION
To address these limitations, the Wuppertal-Budapest lattice QCD collaboration developed a resummation scheme, henceforth referred to as $T^{\prime}$-expansion scheme, that can reach high density with fewer terms in the expansion, exhibits smoother behavior at finite density, and effectively handles the expansion dependence on the QCD transition temperature [19, 20]. This scheme has been used to develop an equation of state with 3D Ising critical point in Ref. [21]. It is, however, crucial to thoroughly assess and explore the performance and limitations of this scheme. To achieve this, we analyze the behavior of the $T^{\prime}$-expansion scheme when applied to effective QCD models designed for two different temperature regimes, namely the Hadron Resonance Gas (HRG) model at low temperatures, and the Cluster Expansion Model (CEM) at high temperatures. A similar analysis of the $T^{\prime}$-expansion scheme was performed recently in Ref. [22] in the context of the functional renormalization group approach.
@@ V Summary
In this work, we investigated the convergence properties of the $T^{\prime}$-expansion scheme, with and without the Stefan-Boltzmann limit correction, compared to the Taylor expansion. The expansion behavior of the scaled baryon density is analyzed in the low-temperature regime using the Hadron Resonance Gas model, and at high temperature using the Cluster Expansion Model. It is observed that the $T^{\prime}$-expansion scheme shows better convergence behavior than Taylor when applied to the Cluster Expansion Model, since this model largely preserves the scaling $\partial/\partial T\sim\partial^{2}/\partial\mu_{B}^{2}$. This is not true for the HRG model: $\kappa_{2}^{B}$ and $\lambda_{2}^{B}$ are monotonically increasing functions in the Ideal case and non-monotonic functions of the temperature for EV- and vdW-HRG models. In all these cases, the $T^{\prime}-$expansion scheme performance is comparable with the one of the Taylor expansion: a similar number of expansion terms is needed, to cover a similar range in $\mu_{B}/T$ with a reasonable agreement with the full result.

### v2

@@ I INTRODUCTION
To address these limitations, the Wuppertal-Budapest lattice QCD collaboration developed a resummation scheme, henceforth referred to as $T^{\prime}$-expansion scheme, that can reach high density with fewer terms in the expansion, exhibits smoother behavior at finite density, and effectively handles the expansion dependence on the QCD transition temperature [19, 20]. This scheme has been used to develop an equation of state with 3D Ising critical point in Ref. [21]. It is, however, crucial to thoroughly assess and explore the performance and limitations of this scheme. This involves comparing different truncation orders to the full result, to check how many terms are needed in each scheme to provide a good description of the full curve, discussing the schemes’ radius of convergence, and comparing the behavior of the (truncated) observables at the same temperatures and chemical potentials, to see e.g. whether the unphysical wiggles that appear in some quantities as a truncation of the Taylor series are also present in the new expansion scheme. To achieve this, we analyze the behavior of the $T^{\prime}$-expansion scheme when applied to effective QCD models designed for two different temperature regimes, namely the Hadron Resonance Gas (HRG) model at low temperatures, and the Cluster Expansion Model (CEM) at high temperatures.
The motivation for choosing these models is the following. The Hadron Resonance Gas is a very common model used to study, among other things, the thermodynamics of strongly interacting matter below the phase transition. Given its simple structure, and given the fact that it is often used to extend lattice QCD thermodynamics to low temperatures, it is important to see how the different expansion schemes work for it. As for the CEM, we chose it because it describes the thermodynamics around the phase transition. There, the Taylor coefficients show a characteristic wiggly structure that is reflected in the Taylor-truncated thermodynamics. It is therefore a good candidate to demonstrate the superiority of the $T^{\prime}-$ expansion scheme. We also tested the expansion scheme on the Holographic model [22] whose higher-order susceptibilities agree with lattice at zero chemical potential but did not include the results since the result is essentially the same as in CEM. A similar analysis of the $T^{\prime}$-expansion scheme was performed recently in Ref. [23] in the context of the functional renormalization group approach.
@@ III.2 Cluster Expansion Model
The CEM uses lattice QCD on leading Fourier coefficients $b_{1}(T)$ and $b_{2}(T)$ of baryon density as input and describes well the resulting high-order baryon number susceptibilities [36]. The CEM provides a lattice-based equation of state model without a critical point. The CEM can be improved by incorporating errors in the lattice inputs. Here it is not essential for the purposes of our studies of convergence properties of the CEM, therefore we do not incorporate lattice errors.
@@ IV.1.2 $T^{\prime}$-expansion scheme
This observation hints that the $T^{\prime}$-expansion scheme behaves as an asymptotic series divergent beyond a certain convergence radius, hence its worsening performance with the inclusion of further terms. We observe that $\mathcal{O}(\mu_{B}^{4})$ seems to provide the optimal performance possible for the HRG class of models.
@@ V Summary
In this work, we investigated the convergence properties of the $T^{\prime}$-expansion scheme, with and without the Stefan-Boltzmann limit correction, compared to the Taylor expansion. The expansion behavior of the scaled baryon density is analyzed in the low-temperature regime using the Hadron Resonance Gas model, and at high temperature using the Cluster Expansion Model and the Holographic model, whose results are not shown as they are similar to those of the Cluster Expansion Model. It is observed that the $T^{\prime}$-expansion scheme shows better convergence behavior than Taylor when applied to the Cluster Expansion Model since this model largely preserves the scaling $\partial/\partial T\sim\partial^{2}/\partial\mu_{B}^{2}$. This is not true for the HRG model: $\kappa_{2}^{B}$ and $\lambda_{2}^{B}$ are monotonically increasing functions in the Ideal case and non-monotonic functions of the temperature for EV- and vdW-HRG models. In all these cases, the $T^{\prime}-$expansion scheme performance is comparable with the one of the Taylor expansion: a similar number of expansion terms is needed, to cover a similar range in $\mu_{B}/T$ with a reasonable agreement with the full result. We also observe that while $T^{\prime}$ expansion offers similar or superior performance to Taylor expansion at low orders, its advantages at higher orders appear to be limited. Therefore, there is no evidence from model studies that evaluating high-order $T^{\prime}$-expansion coefficients, which is very expensive to do with lattice QCD, will offer substantial advantages in describing the equation of state at finite baryon density.